Diana inherits $45,000 and decides to invest part of it in an education account for her daughter and the rest in a 10-year CD. If the amount she puts in the education account is $12,000 more than twice the amount she puts in the CD, how much money does Diana invest in each account? Define unknown quantities with one variable and show all work.
Found 2 solutions by ankor@dixie-net.com, addingup:
Answer by ankor@dixie-net.com(22740)

Diana inherits $45,000 and decides to invest part of it in an education account for her daughter and the rest in a 10-year CD.
If the amount she puts in the education account is $12,000 more than twice the amount she puts in the CD, how much money does Diana invest in each account?
:
let x = amt invested in the CD
"the amount she puts in the education account is $12,000 more than twice the amount she puts in the CD,", therefore
2x + 12000 = amt invested in the ed acct
:
cd + ed acct = $45,000
x + (2x+12000) = 45000
3x + 12000 = 45000
3x = 45000 - 12000
3x = 33000
x = 33000/3
x = $11,000 invested in the cd
then
2(11000) + 12000 = $34,000 invested in the ed acct

Answer by addingup(3677)

Let the amount invested in the CD be x:
x+2x+12000 = 45000
3x+12000 = 45000
3x = 33000
x = 11000 this is the CD
The education acct:
2*11000+12000 = 34000